OpenStudy (anonymous):

You would like to join a fitness center. Fit-N-Trim charges $80 per month. Fit-For-Life charges a one time membership fee of $75 and $55 per month. Which fitness center offers the better rate? Explain.

OpenStudy (anonymous):

Lets say you want to join a fitness club for 12 months. Lets compare the price... (a) Fit - N - Trim : 80 per month for 12 months = (80)(12) = 960 (b) Fit - for - Life : 75 + 55(12) = 75 + 660 = 735 The better deal here is (b)Fit - For - Life. Now lets see what it would be if you only wanted to join a fitness club for 2 months. (a) Fit - N - Trim : 80 per month for 2 months = (80)(2) = 160 (b) Fit - For - Life : 75 + 55(2) = 75 + 110 = 185 The better deal here is (a) Fit - N - Trim. The only time they would be equal would be : 80x = 75 + 55x .....x being the number of months 80x - 55x = 75 25x = 75 x = 75/25 x = 3 The only time they would cost the same is if you only went for 3 months So, as to your question, it depends on how long you wanted to go for. Anything less then 3 months, the better deal would be Fit - N - Trim. Anything more then 3 months, the better deal would be Fit - For - Life. If you only went for 3 months, it wouldn't matter because they both cost the same.